The concentrated charges of 0.25microcoulombs each are located at the vertices of an equilateral triangle of 10cm side. Find the magnitude and direction of force on one charge due to other charges


1
Expert's answer
2021-11-10T10:10:45-0500

"F_1=F_2=\frac{kq^2}{r^2}=0.056~\text N,"

"F=F_1\sqrt 3=0.1~\text N."